Compare the shift in frequency if a 2000Hz source is moving towards you at 15ms-1 versus if you are moving towards it at 15ms-1. Are the two frequencies exactly the same? Are they close? Repeat

the calculation for 150ms-1 and then again for 300ms-1. What can you conclude about the asymmetry of the Doppler formulas? Show that at low speeds (relative to the speed of sound), the two formulas-source approaching and detector approaching-yield the same result.



1
Expert's answer
2021-03-09T15:31:08-0500

"v_{snd} = 343 \;m/s"

At 15 m/s

"f’_{source} = f\frac{1}{1 -\frac{v_{stc}}{v_{snd}}} \\

= 2000 \times \frac{1}{1 -\frac{15}{343}} \\

= 2092 \;Hz \\

f’_{observer} = f(1 + \frac{v_{stc}}{v_{snd}}) \\
=2000(1 +\frac{15}{343}) \\
= 2087 \;Hz"

At 150 m/s

"f’_{source} = f\frac{1}{1 -\frac{v_{stc}}{v_{snd}}} \\

= 2000 \times \frac{1}{1 -\frac{150}{343}} \\

= 3554 \;Hz \\

f’_{observer} = f(1 + \frac{v_{stc}}{v_{snd}}) \\
=2000 (1 +\frac{150}{343}) \\
= 2874 \;Hz"

At 300 m/s

"f’_{source} = f\frac{1}{1 -\frac{v_{stc}}{v_{snd}}} \\

= 2000 \times \frac{1}{1 -\frac{300}{343}} \\

= 15,961 \;Hz \\

f’_{observer} = f\frac{1}{1 + \frac{v_{stc}}{v_{snd}}} \\
=2000 (1 +\frac{300}{343}) \\
= 3749 \;Hz"

At low speeds (relative to the speed of sound), the two formulas-source approaching and detector approaching-yield the same result.

"f’_{source} = f\frac{1}{1 -\frac{v_{stc}}{v_{snd}}} \\

f’_{source} = f(1 -\frac{v_{stc}}{v_{snd}})^{-1} \\

f’_{source} ≈ f(1 +\frac{v_{stc}}{v_{snd}}) \\

f’_{source} = f’_{observer}"